Please show the process on how to complete this.After researching the competitors of EJH Enterprises, you determine that most comparable firms have the following valuation r
Answer #1

As a first step let's calculate the minimum and maximum of each of the two multiples given. Please see the last two columns in the table below.

Comp 1 Comp 2 Comp 3 Comp 4 Min Max
EV/ EBITDA             12.00             11.00     12.50        10.00             10.00          12.50
P/E             19.00             18.00     20.00        17.00             17.00          20.00

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics.

Parameter Linkage Min Max
EV/ EBITDA multiple average A              10.00                 12.50
EBITDA of EJH B           290.00              290.00
EV C = A x B        2,900.00           3,625.00
[-] Debt D              41.00                 41.00
[+] Cash E              29.00                 29.00
Equity value F = C - D - E        2,888.00           3,613.00
Nos of shares outstanding N           102.00              102.00
Hence, price per share G = F / N              28.31                 35.42
P/E Multiple average H              17.00                 20.00
EPS of EJH I                1.80                   1.80
Price per share J = H x I              30.60                 36.00

Hence, lowest price with both ranges = max (Min value with EV/EBITDa method, Min value with P/E method) = max (28.31, 330.60) = $ 30.60

And, highest price with both ranges = min (Max value with EV/EBITDa method, Max value with P/E method) = min (35.42, 36.00) = $ 35.42